OpenStudy (anonymous):

What is the exact value of the expression sqrt72 − sqrt8 +sqrt128 ? Simplify if possible.

jhonyy9 (jhonyy9):

72=9*4*2 so than sqrt72=3*2sqrt2=6sqrt2 sqrt8=2sqrt2 sqrt128=sqrt(64*2)=8sqrt2 so than 6sqrt2 -2sqrt2 +8sqrt2=12sqrt2

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

aweso0me :) thank you

jhonyy9 (jhonyy9):

merry christmas ! good luck bye

OpenStudy (anonymous):

another way of showing your answer: